- Unit Log -
|SS: early-70s perhaps (guess based on componentry),
tube: 1952 (3.5MHz crystal stamped with "6-52") as Type JX-10
|Date of Receipt:||2016 (for repair & reimplementation)|
|State upon Receipt:||Non-working, presumed due to failed negative switching supply, poor workmanship in conversion.|
Main chassis stripped and cleaned.
Covers made for grosser holes in chassis.
IEC mains connector installed.
IF strip constructed and wired in.
Power supply and audio board constructed and installed.
Detector board prepared and mounted.
|Date:||2018 Fall - 2019 Spring|
Front panel wired up.
Detector, AVC, Meter circuitry installed.
Lo-band buffer, 2nd conversion, lo-hi band switching constructed and installed.
BFO constructed and installed. Coil section rewound (50 turns * 2 strands #36 wire).
BFO Oscillator modifications:
Source R added for self-bias. Load choke replaced with R. Coupling C reduced to 68pF to minimise influence from raised base of diff amp. 2SK19 used instead of MPF102 but also works with 102.
BFO Mixer modifications:
Emitter R lowered to 750 to raise output level. Coupling C from IF reduced to 10pF. IF buffer not needed.
|Procedure:||100KHz calibrator circuit inverted to allow use of NPN transistor. Power isolation resistor and cap installed inside module, zener removed.|
|Procedure:||Front end electronics strip with tuning cap removed, HFO buffer attached with cap, E resistor value increased, functional now for feeding a frequency counter.|
|Procedure:||Attempt to reduce spurious oscillations on high bands. Tuning on L33/34 also was not very sharp. L33/L34 mica caps replaced, 51pF for 51pF, 68pF for 65pF. Additional 0.001 bypass cap added inside can.|
|Procedure:||Band D (7.4-14.5) F aligned with frequency counter and WWV and peaked at lo/hi ends. Frequency accuracy improved. Alignment of signal sections vague due to lack of signal generator.|
|Procedure:||Band A (0.54-1.4) F aligned with frequency counter and peaked at lo/hi ends.|
|Procedure:||Incandescent dial lamps replaced with bayonet-socket LEDs. Power for lamps moved to 5V.|
|Procedure:||Unit returned to owner.|